OpenStudy (anonymous):

HELLPPPP All of the following equations have the same solution except _____. 7.5x= 1.5 = 0.04 x+ 6.5 = 6.3 -1.8 +x= -1.6

OpenStudy (anonymous):

B is x/5 = 0.04

OpenStudy (anonymous):

C is -0.2 and the rest is 0.2

OpenStudy (anonymous):

so its c?

OpenStudy (anonymous):

yeah. but do you know how to do this or do you want me to show you?

OpenStudy (anonymous):

show me

OpenStudy (anonymous):

you must get the x alone like this a: 7.5x =1.5 x=1.5/7.5 x=0.2 b: x/5=0.04 (x*5)/5 =0.04*5 x=0.2 c: x+6.5=6.3 x=6.3-6.5 x=-0.2 d: -1.8+x=-1.6 x=-1.6+1.8 x=0.2